OpenStudy (anonymous):

Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)? A. y = x + 1 B. y = x + 1 C. y = x – 1 D. y = x – 1 i got D i just want to make sure this is correct?

jimthompson5910 (jim_thompson5910):

It's not D, here's why y = x - 1 -1 = -2 - 1 ... Plug in x = -2 and y = -1 -1 = -3 which is false So (-2,-1) does NOT lie on the line y = x-1

jimthompson5910 (jim_thompson5910):

btw, A and B are the same...and C and D are the same

OpenStudy (anonymous):

i am so sorry i typed it in wrong A. Y=1/2x+1 B.Y=x+1 C.Y=1/2x-1 D.Y=x-1

jimthompson5910 (jim_thompson5910):

that's ok, but it's still not D for the same reason explained above

OpenStudy (anonymous):

alright thank you for that explanation :)

jimthompson5910 (jim_thompson5910):

np, what else did you get

OpenStudy (anonymous):

i got the answer B

OpenStudy (anonymous):

i did -1=-2+1 -1=-1

jimthompson5910 (jim_thompson5910):

choice B is correct

OpenStudy (anonymous):

but what about the (-4,-3)?

jimthompson5910 (jim_thompson5910):

y = x + 1 -3 = -4 + 1 -3 = -3 which is true

jimthompson5910 (jim_thompson5910):

so (-4,-3) lies on y = x+1

OpenStudy (anonymous):

ohh it does go in! thank you sir for all the help :)

jimthompson5910 (jim_thompson5910):

you're welcome